OpenStudy (anonymous):

A swimmer goes from A to B in 4 hr and return back from B to A in only 2 hr. find velocity of river flow and velocity of swimmer in still water.

OpenStudy (anonymous):

Let the velocity of the river be \(v_r\) and the velocity of the swimmer be \(v_s\). Let the time to to go the first trip be \(\Delta t_{AB}\) and the time for the reverse trip be \(\Delta t_{BA}\). Let the distance between point \(A\) and \(B\) be \(\ell\). By the definition of constant velocity, we can write the two following equations that you can use to solve for the velocities.\[v_s-v_r=\frac{\ell}{\Delta t_{AB}}\]\[v_s+v_r=\frac{\ell}{\Delta t_{BA}}\]

OpenStudy (anonymous):

i think it go over my head.can u xplain it more clearly/.

OpenStudy (anonymous):

Oops, I forgot the distance, x-> speed of the man in still water y-> speed of the river current \(x+y =\frac l 4\) \(x-y =\frac l 2 \)

OpenStudy (anonymous):

@Ruchi. What is confusing you about my post?

OpenStudy (anonymous):

i hav nt understood it clearly what ever u hav written @yakeyglee

OpenStudy (anonymous):

What trips you up?